OpenStudy (nuttyliaczar):

Find the electric field a distance z above the center of a square loop (with side length a) carrying a uniform line charge. I mostly need help with setting up the integral here. The tricky part is changing the unit vector r^ to something I can integrate. This is a physics question but I figured that asking on here isn't too ridiculous.

OpenStudy (nuttyliaczar):



OpenStudy (nuttyliaczar):

This gives a visual of what the problem is. R is the shortest length r can be (thus it is a constant). x is the base of the arbitrary triangle.

OpenStudy (nuttyliaczar):

\[E=\frac{ 1 }{ 4\pi \epsilon }\int\limits_{y}^{z}\frac{ \lambda }{ r ^{2} }r _{v}dr\]

OpenStudy (nuttyliaczar):

This is the equation for the electric field given some distance r for anyone who doesn't know. \[r _{v}\] is the unit vector of r.

OpenStudy (nuttyliaczar):

Can anyone explain this more thoroughly to me? The person asking the question in that link seems to have done most of the problem already, but I don't know how he got it or how much of it is correct.

OpenStudy (anonymous):

the solve from link is correct

OpenStudy (anonymous):

@nuttyliaczar Hi, I've found a good solution for it. Do you still like to discuss this problem?

OpenStudy (anonymous):

Did you find a way to make your integral?

OpenStudy (anonymous):

As this question was very beautiful and i liked it so much I tried to solve it. Making the integral wasn't that hard. The hard problem for me was calculating it. I asked for help and guidance on calculations here: http://openstudy.com/study#/updates/567b4085e4b0f8b8229fbf99

OpenStudy (anonymous):

after a very long time, i solved that hard integral with some big help. But let me know what you found when solving it.

OpenStudy (anonymous):

@nuttyliaczar seems u're not around. the integral can be written as below: \[\int\limits dE=\frac{8 k.q }{ l } \int\limits_{0}^{\frac{ a }{ 2 }}[(x-\frac{ a }{ 2 })^2+B]^{-\frac{ 3 }{ 2 }}dx\] Where: k=9*10^9 q=total static charge l= total length of the frame \[B=z^{2}+\frac{a ^{2} }{ 4 }\]

OpenStudy (anonymous):

its limits are x=0 and x=a/2 The Intergral above gives the total E at the point P. We can talk about it later if you still like to know how this can be derived.